itsdakdidd:

Patel squeezed oranges so that his family could have fresh-squeezed juice for breakfast. He squeezed StartFraction 4 over 17 EndFraction cups from the first orange, StartFraction 3 over 10 EndFraction cups from the second orange, StartFraction 9 over 20 EndFraction cups from the third orange, StartFraction 3 over 11 EndFraction cups from the fourth orange, and StartFraction 7 over 15 EndFraction cups from the fifth orange. Patel estimates that he needs 3 cups of orange juice for his family. About how much more orange juice does he need to reach his estimate?

mhanifa:

Can you sum the fractions?

mhanifa:

\[\frac{ 4 }{ 17 }+\frac{ 3 }{ 10 }+\frac{ 9 }{ 20}+\frac{ 3 }{ 11 }+\frac{ 7 }{ 15 }+x = 3\] Solve the equation for x.